Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
seglararro, derossi, amagazzu, gherardini - 0:25 Saturday 30 March 2024 (63794) Print this report
SRCL set scan with hrec/hinj and optical response measurement
The goal of the shift was to have various measurements of the error on h(t) reconstruction at low frequencies, as a follow-up investigation to what was reported in logbook entry 63519. The SRCL set point was changed smoothly and fixed at 7 values (Fig1). At each SRCL set point, we perform line injections in a good set of frequencies, to obtain hrec/hinj (same lines as checkhrec measurements) and broadband injections to have the optical response of NE, WE and BS.

We started at 18:57 UTC and finished at 23:01 UTC, due to an unlock.

- First unlock happened at 19h24 from LN3 due to PR_F0_TY (Fig2)
- Second unlock from CARM_NULL_1f due to DIFFp_TY oscillation at 4Hz (Fig3)
- The cause of the third unlock from LN3 is unclear but seems like there is a glitch in NE_MIR_zCorr (Fig4)

Images attached to this report
Comments to this report:
verkindt - 13:40 Saturday 30 March 2024 (63795) Print this report

First plot shows the spectrogram of hoft_raw, where can be seen the checkhrec injections (lines injections during a few minutes) and optical response injections (broadband injections over about 10mn). Second plot shows the change of SRCL_SET value and the changes of the modulus and phase of hrec/hinj estimated by Hrec for a few permanent calibration lines. The large fluctuations of hrec/hinj modulus and phase are when the check or optical response injections have been done. In between, we can see that hrec/hinj modulus and phase are affected by the change of SRCL_SET (tuning of SR) mainly at low frequencies. This should be confirmed by the checkhrec injections results to come soon.

Images attached to this comment
seglararro - 12:59 Friday 05 April 2024 (63852) Print this report
We analysed the data taking during the shift. For simplicity, the several steps with numbers, as well as the OS (sampled at 0.01Hz and 0.001Hz) can be seen in Figure 1. The results of the hrec/hinj computation for injections 1-2-3-4-6-7 are attached measured by the NE EM actuator, WE EM actuators, WE PCal and NE Pcal, in Figures 2,3,4,5, respectively. Note that for the Pcal, two measurements are obtained per SRCL_SET step. The averages hrec/hinj are shown in Figures 6,7,8,9. We computed the corresponding optical responses at each of these steps (1-2-3-4-5-6), which can be found in Figure 10 for NE and Figure 11 for WE.

In these measurements, we did not find a flat response at low frequencies. Yet, the 30th (a day after), during daily calibration injections, we see a ~flat response for NE and WE (WE in Figure 12), for SRCL_SET_TOT~35 and OS_LF_HZ (first green line and black line used to computed the OR, Figure 13).

Note that these results use updated Pcal models, and EM actuators models not yet updated to the latest measurements (as these are currently under study).
Images attached to this comment
Search Help